OpenStudy (anonymous):

A teacher wants to send 4 students to the library. If there are 21 students in the class, how many ways can he choose to send 4 students

OpenStudy (kropot72):

Th required number of combinations is given by \[21C4=\frac{21!}{4!17!}=\frac{21\times 20\times19 \times18}{4\times 3\times2 \times1}=you\ can\ calculate\]

OpenStudy (anonymous):

thank you!